Find the length of the unknown side of each right triangle with legs a and b and hypotenuse c. Round to three decimal places, if necessary.
1) a = 3, b = 9

A) c = 27

B) c = 9.487

C) c = 12

D) c = 4.899

Answer by stanbon(48546)

c^2= 3^2+9^2
c^2=9+81
c^2=90
c=3sqrt10
c=9.4868....
Cheers,
Stan H.